OpenStudy (anonymous):

Solve x ^2 - 8x + 3 = 0 by completing the square. Which equation is used in the process? (x - 4)^2 = 0 (x - 4)^2 = 13 (x - 4)^2 = 19

OpenStudy (johnweldon1993):

First we must get this in the correct form \[\large x^2 - 8x = -3\] To complete the square, we take the coefficient of the 'x' ....divide it by 2...then square the result. So here, -8 would be the coefficient Divide that by 2...so -4....and square the result (-4)^2 = 16 So now we add this to both sides of that equation \[\large x^2 - 8x + 16 = -3 + 16\] \[\large x^2 - 8x + 16 = 13\] Write the left hand side as a square \[\large (x - 4)^2 = 13\]

OpenStudy (anonymous):

got it thanks you make it look easy

OpenStudy (johnweldon1993):

Not a problem! And everything in math is easy once you practice it enough!